# SEDMamba: Enhancing Selective State Space Modelling with Bottleneck Mechanism and Fine-to-Coarse Temporal Fusion for Efficient Error Detection in Robot-Assisted Surgery

### v1

@@ Abstract
Automated detection of surgical errors can improve robotic-assisted surgery. Despite promising progress, existing methods still face challenges in capturing rich temporal context to establish long-term dependencies while maintaining computational efficiency. In this paper, we propose a novel hierarchical model named SEDMamba, which incorporates the selective state space model (SSM) into surgical error detection, facilitating efficient long sequence modelling with linear complexity. SEDMamba enhances selective SSM with bottleneck mechanism and fine-to-coarse temporal fusion (FCTF) to detect and temporally localize surgical errors in long videos. The bottleneck mechanism compresses and restores features within their spatial dimension, thereby reducing computational complexity. FCTF utilizes multiple dilated 1D convolutional layers to merge temporal information across diverse scale ranges, accommodating errors of varying durations. Besides, we deploy an established observational clinical human reliability assessment tool (OCHRA) to annotate the errors of suturing tasks in an open-source radical prostatectomy dataset (SAR-RARP50), constructing the first frame-level in-vivo surgical error detection dataset to support error detection in real-world scenarios. Experimental results demonstrate that our SEDMamba outperforms state-of-the-art methods with at least 1.82% AUC and 3.80% AP performance gain with significantly reduced computational complexity. [^1] The corresponding dataset, code and models will be released upon paper acceptance.
@@ I Introduction
Observational clinical human reliability assessment (OCHRA) is a standardized method to assess the quality of surgical execution by detecting and characterising the nature (procedural or executional) and severity of technical errors on various specialties [4, 5, 6]. Clinical validation has identified executional errors, following the OCRA definitions, as a strong predictor of both skill and patient outcomes [7, 8]. This strongly promotes the application of such tools for reducing adverse events and surgical training. Deployment of OCHRA is impractical due to the significant time required for reviewing and manual annotation. Subjectivity in validity and reliability issues further prioritize the need for automated solutions [9, 10].
@@
Propelled by deep learning (DL) methods and the public JIGSAWS dataset [11], containing atomic gesture and skill annotations from RAS tasks in dry labs, significant prior work on RAS video understanding for skill analysis focused on recognition and temporal localization of surgical gestures [12, 13, 14, 15, 16]. Recent work adapted DL methods for gesture recognition to in-vivo datasets[17, 18, 19], and also studied the association of the observed gesture sequence to patient outcomes and surgeon’s skill [20]. Detection and temporal localization of procedural and executional errors have not been extensively studied. Hutchinson et al. [21] incorporated error annotations in two JIGSAWS tasks (simulation suturing and needle passing). Procedural errors are considered when a surgical gesture is omitted or the task is carried out with a gesture sequence that deviates from the defined grammar graph, while executional errors denote moments of suboptimal performance. For each gesture, a maximum of four errors were considered from a list (multiple attempts, needle orientation, needle drop, out of view), used to characterize the entire (begin to end) gesture as erroneous (if any error was noted) or normal [21]. Employing these annotations, Li et. al. compared three DL classifiers for detecting erroneous gestures [22].
We argue that the annotation approach in [21], though successful in dry-lab, is not directly applicable to in-vivo RAS settings. It is not possible to consider the same error definitions for all gestures. It is also typical for errors to span across multiple gestures, thus their temporal localization can not always align to the temporal boundaries (begin, end) of gestures. Consequently, its annotation based on pre-segmented gesture clips fails to accurately provide temporal boundaries for the start and end of errors. In this work, we define custom error descriptions, following OCHRA, and present the first in-vivo RAS dataset with frame-level error annotations to provide more accurate temporal boundaries of errors for surgical error detection (SED).
@@
SED aims to detect errors that occur during surgery. In [22], kinematic data was further downsampled to lengths of 30 using sliding windows within manually pre-segmented gesture clips, serving as the input for a CNN-LSTM network to detect executional errors. However, its parallel computing capability is constrained by long short-term memory (LSTM). Samuel and Cuzzolin [23] developed a deep convolutional autoencoder to detect errors in pre-trimmed video clips in an unsupervised manner, yet the local receptive field of convolution layers hinders its global temporal modelling capability. Overall, existing research on SED has primarily focused on short kinematics data and video clips, efficient SED in long surgical videos remains unexplored. To achieve this objective, it is essential to consider two key factors: long-term dependency modelling and computational complexity. Transformer architectures, benefiting from the self-attention mechanism, enable global modelling and have demonstrated impressive performance in surgical video analysis tasks such as surgical phase recognition [24, 25]. Nevertheless, the quadratic computational complexity of self-attention with respect to video sequence length significantly increases the computational burden, impeding its efficiency in handling long video scenarios.
@@
Figure 1: A surgical video from the SAR-RARP50 dataset, including errors of varying durations.
@@
Recently, state space models (SSMs) have proven highly effective in capturing long sequences in the natural language processing domain [26, 27]. Selective SSM [27], namely Mamba, distinguishes itself from other SSM-based methods by integrating a data-dependent SSM with a selection mechanism that employs parallel scanning, thereby achieving linear complexity while preserving global receptive fields. Inspired by the great success of Mamba, researchers have explored it in the computer vision domain, yielding promising performance. Zhu et al. [28] proposed a bidirectional Mamba (Vim) that scans images forward and backward to enhance spatial-aware understanding for image tasks, significantly improving computational efficiency while achieving higher performance compared to well-established Transformers. SegMamba [29] constructed a tri-oriented Mamba for 3D medical image segmentation. However, as depicted in Fig. 1, surgical videos contain errors with varied durations, necessitating the algorithm to consider temporal information at different granularities. Prior selective SSM-based works rarely consider modelling multi-scale temporal information, we thus showcase the effectiveness of a novel selective SSM-based method named SEDMamba, which captures temporal information across various scales to detect surgical errors in suturing from robotic-assisted radical prostatectomy (RARP) cases.
@@
    We propose a novel hierarchical model, SEDMamba, constructed with bottleneck multi-scale state space blocks to detect surgical errors in long videos. To handle errors with variable durations, we design a fine-to-coarse temporal fusion (FCTF) using different dilated 1D convolutional layers to merge temporal information across different scale ranges. Besides, to further reduce computational complexity, we incorporate a bottleneck mechanism that compresses the spatial dimension of features, enabling selective SSM to concentrate on modelling the temporal dimension. To the best of our knowledge, this is the first attempt to exploit selective SSM in surgical error detection.
@@
    We present the first frame-level surgical error annotations of an in-vivo RAS dataset.
@@
    The proposed SEDMamba outperforms state-of-the-art video analysis methods and selective SSM-based methods by 1.82-13.06% AUC and 3.8-18.86% AP while requiring 11.71-75.79% fewer model parameters and 40.71-90.93% fewer FLOPs.
@@ II-A Dataset Description and Annotation
A total of 48 videos with 60 Hz from the open-source, SAR-RARP50 [30, 31], dataset of the RARP cases were independently and blindly reviewed by two assessors. SAR-RARP50 focuses on the suturing of the dorsal venous complex (DVC) task and contains open-source annotations of eight gestures used as a task analysis sheet [30]. We developed a customized tool for the annotation of both procedural and executional errors using a previously validated suturing checklist and the OCHRA methodology [32]. This was summarized as a list of possible procedural and executional errors within any given suturing task, as seen in Table I. The duration of these videos varies from one minute to over eleven minutes, including the complete suturing process. Two assessors (one senior and one junior RAS surgeon with expertise in RAS education) analyzed these untrimmed videos sequentially frame by frame and annotated the start and end frame indexes of all errors. Fig. 2 illustrates three types of errors from our annotations. Inter-rater reliability was verified with Pearson’s correlation showing very strong agreement ($r$=0.8726, $p<$0.001) between the two assessors. Finally, consensus was established for all disagreements via joint review.
@@
Multiple attempts (end of 1st attempt
@@
E7 Incorrect position along needle (normally 2/3rds) 89 1737
@@
A total of 2507 error cases were annotated, with E24-“Incorrect/poor instrument control” occurring the most frequently (686), followed by E1-“Multiple attempts” (510). The frequency of error types had statistically significant difference ($\chi 2$ = 1945.15, $p<$ 2.2e-16). The duration of errors varies, with the longest average duration being E19-“Suture entanglement”, averaging around 599 frames per case, while the shortest is E11-“Grasped at needle tip”, averaging 8 frames per case. The 48 videos comprise a total of 976037 frames. In the SED task, the frame is categorized as “normal” if no error exists or “error” if any type of error exists, resulting in 320355 error frames. To detect errors of varying durations, irrespective of their types, it is particularly crucial to model temporal information across multiple scales.
@@ II-B Surgical Error Detection Model (SEDMamba)
This section introduces the preliminary concepts related to SEDMamba, including state space modelling and discretization. We then formulate the SEDMamba architecture and its core bottleneck multi-scale state space (BMSS) block with fine-to-coarse temporal fusion.
@@ II-B2 Overview of SEDMamba
SEDMamba has a hierarchical architecture, shown in Fig. 3, comprising three bottleneck multi-scale state space (BMSS) blocks. Taking a video $V\in\mathbb{R}^{L\times C\times H\times W}$ with $L$ length, we first extract the spatial embedding sequence $\mathcal{E}\in\mathbb{R}^{L\times D}$ from a fixed DINOv2 [33], where $D$ is the spatial dimension. DINOv2 is a strong vision foundation model proven to extract task-agnostic visual spatial embedding and excel in video-based tasks [33]. Then, SEDMamba compresses $\mathcal{E}$ in the spatial dimension and captures fine-to-coarse temporal feature through each BMSS block, resulting in $i$-th hierarchical spatial-temporal representations with size $L\times\frac{D}{2^{i}}$. The 1D convolutional layer with 1 kernel size and 1 channel is then used for binary classification and outputs the final error probability prediction $P\in\mathbb{R}^{L\times 1}$. Note that each value in $P$ indicates the likelihood that the corresponding frame is erroneous. We use binary cross-entropy loss $\mathcal{L}_{bce}$ for supervised training.
@@ II-B3 Bottleneck Multi-scale State Space (BMSS) Block
20:                  $\left.\left.\text{Broadcast}^{G}(\text{Linear}_{\Delta}^{1}(x))\right)\right)$
@@
II-B4 Fine-to-Coarse Temporal Fusion (FCTF)
To tackle varying durations of error and restricted receptive fields of convolutional layers, we propose the FCTF to capture and merge multi-scale temporal information. FCTF stacks three dilated 1D convolution layers with dilation rates of 2, 4, and 8 from top to bottom, enabling the fine-to-coarse temporal feature extraction scale by scale. Besides, stacking dilated convolution increases the receptive field without the need to increase the parameter number by increasing the kernel size. The receptive field expands exponentially as the number of stacked layers increases. Therefore, with a few parameters, we achieve a significantly large receptive field in the temporal dimension, which mitigates model overfitting and facilitates efficient error detection. The receptive field at each layer, with a kernel size of 3, can be calculated with the following formula:
$\text{ReceptiveField}(l)=2^{l+2}-1$ (5)
where $l\in[1,3]$ represents the layer number.
@@ III-A Experimental Settings
All experiments are implemented in PyTorch on a GTX 4090Ti GPU. We use AdamW optimizer with initial learning rate of 1e-4 and maximum training epoch set at 200. We sampled SAR-RARP videos at 5Hz to get 40 videos with 46,586 normal frames and 18,540 error frames as the training set, and 8 videos with 7549 normal frames and 2096 error frames as the testing set. We use DINOv2 with fixed weights pretrained on ImageNet [33] as the visual extractor to generate spatial embedding sequences $\mathcal{E}$ from the video with $L$ frames as the model input. The compression factor $G$ in the BMSS block is set as 64. The channel $E$ of dilated 1D convolutional layers in FCTF is fixed as $\frac{D}{8}$. Following  [27], the state dimension $N$ of selective SSM is set to $16$. We the standard and widely used evaluation metrics: area under the curve (AUC) and average precision (AP) for the SED task. We report the mean and standard deviation from three runs with different random seeds.
III-B Comparison with the State-of-the-Art
We compare the proposed SEDMamba with state-of-the-art convolutional and Transformer-based video analysis methods: TeCNO [35], MS-TCN [36], MS-TCN++ [37], ASFormer [24], and selective SSM-based methods: Mamba [27] and Vim [28]. All methods are implemented based on their released code and original literature, and fine-tuned to fit our error detection task to provide a basis for a fair comparison.
Table II shows quantitative results of all methods. The proposed SEDMamba consistently achieves superior performance with an increase of 1.82-13.06% in AUC and 3.8-18.86% in AP, compared to other methods. Notably, all selective SSM-based methods surpass the Transformer-based method ASFormer, demonstrating the superiority of selective SSM in the SED task. Fig. 4 illustrates error detection results from all test videos. The second-best method Mamba tends to output smoother and more ambiguous prediction curves, which impedes its sensitivity to errors. It also produces lower confidence outputs within error frames. Thanks to the fine-to-coarse temporal information provided by FCTF for selective SSM, SEDMamba is able to identify suddenly occurring errors with short duration (highlighted in red boxes) and consistently produces high confidence within error areas in most cases (highlighted in black boxes). In most normal frames, SEDMamba has comparable or lower error probabilities to Mamba, with only a small portion showing higher error probabilities. Note that both AUC and AP comprehensively consider all classes, i.e. error and normal, indicating SEDMamba has superior overall performance.
TABLE II: Quantitative results of SEDMamba and other methods. Higher AUC and AP mean better performance. Top two results are highlighted in bold and underlined.
Methods AUC (%) $\uparrow$ AP (%) $\uparrow$
TeCNO [35] 58.14$\pm$0.33 26.01$\pm$0.32
MS-TCN [36] 58.82$\pm$0.16 29.98$\pm$3.06
MS-TCN++ [37] 58.57$\pm$0.43 30.35$\pm$1.35
ASFormer [24] 68.76$\pm$0.93 36.79$\pm$2.35
Mamba [27] 69.38$\pm$0.26 41.07$\pm$0.95
Vim [28] 69.06$\pm$1.46 40.63$\pm$2.30
SEDMamba (ours) 71.20$\pm$0.26 44.87$\pm$1.52

### v2

@@ Abstract
Automated detection of surgical errors can improve robotic-assisted surgery. Despite promising progress, existing methods still face challenges in capturing rich temporal context to establish long-term dependencies while maintaining computational efficiency. In this paper, we propose a novel hierarchical model named SEDMamba, which incorporates the selective state space model (SSM) into surgical error detection, facilitating efficient long sequence modelling with linear complexity. SEDMamba enhances selective SSM with a bottleneck mechanism and fine-to-coarse temporal fusion (FCTF) to detect and temporally localize surgical errors in long videos. The bottleneck mechanism compresses and restores features within their spatial dimension, thereby reducing computational complexity. FCTF utilizes multiple dilated 1D convolutional layers to merge temporal information across diverse scale ranges, accommodating errors of varying duration. Our work also contributes the first-of-its-kind, frame-level, in-vivo surgical error dataset to support error detection in real surgical cases. Specifically, we deploy the clinically validated observational clinical human reliability assessment tool (OCHRA) to annotate the errors during suturing tasks in an open-source radical prostatectomy dataset (SAR-RARP50). Experimental results demonstrate that our SEDMamba outperforms state-of-the-art methods with at least 1.82% AUC and 3.80% AP performance gains with significantly reduced computational complexity. The corresponding error annotations, code and models will be released at https://github.com/wzjialang/SEDMamba.
@@ I Introduction
Observational clinical human reliability assessment (OCHRA) is a standardized method to assess the quality of surgical execution by detecting and characterising the nature (procedural or executional) and severity of technical errors on various specialties [4, 5, 6]. Clinical validation has identified executional errors, following the OCRHA definitions, as a strong predictor of both skill and patient outcomes [7, 8]. This strongly promotes the application of such tools for reducing adverse events and surgical training. Deployment of OCHRA is impractical due to the significant time required for reviewing and manual annotation. Subjectivity in validity and reliability issues further prioritize the need for automated solutions [9, 10].
@@
Propelled by deep learning (DL) methods and the public JIGSAWS dataset [11], containing atomic gesture and skill annotations from RAS tasks in dry labs, significant prior work on RAS video understanding for skill analysis focused on recognition and temporal localization of surgical gestures [12, 13, 14, 15, 16]. Recent work adapted DL methods for gesture recognition to in-vivo datasets[17, 18, 19], and also studied the association of the observed gesture sequence to patient outcomes and surgeon’s skill [20]. Detection and temporal localization of procedural and executional errors have not been extensively studied. Hutchinson et al. [21] incorporated error annotations in two JIGSAWS tasks (simulation suturing and needle passing). Procedural errors are considered when a surgical gesture is omitted or the task is carried out with a gesture sequence that deviates from the defined grammar graph, while executional errors denote moments of suboptimal performance. For each gesture, a maximum of four errors were considered from a list (multiple attempts, needle orientation, needle drop, out of view), used to characterize the entire (begin to end) gesture as erroneous (if any error was noted) or normal [21]. Employing these annotations, Li et. al. compared four DL classifiers for detecting erroneous gestures [22]. We argue that the annotation approach in [21], though successful in dry-lab, is not directly applicable to in-vivo RAS settings. It is typical for errors to span across multiple gestures, thus their temporal localization can not always align to the temporal boundaries (begin, end) of gestures. Consequently, its annotation based on pre-segmented gesture clips fails to accurately provide temporal boundaries for the start and end of errors. In this work, we define custom error descriptions, following OCHRA, and present a novel in-vivo RAS dataset with first-of-its-kind, frame-level error annotations to provide more accurate temporal boundaries of errors for surgical error detection (SED).
@@
SED aims to detect errors that occur during surgery. It has been formulated as a binary classification task, where the ground truth labels are ‘normal’: if no error exists, or ‘error’: if any error type exists [22, 23]. In [22], kinematic data was downsampled to segments of 30 data points, using sliding windows within manually pre-segmented gesture clips, serving as the input for a long short-term memory (LSTM) network to classify each gesture clip as either error or normal. However, its parallel computing capability is constrained by the inherently state-dependency nature of LSTM. Samuel and Cuzzolin [23] developed an unsupervised convolutional autoencoder to categorise each frame as either error or normal in pre-trimmed video clips, yet the local receptive field of convolution layers hinders its global temporal modelling capability. Overall, existing research on SED has primarily focused on short kinematics data and video clips, efficient SED in long surgical videos remains unexplored. To achieve this objective, it is essential to consider two key factors: long-term dependency modelling and computational complexity. Transformer architectures, benefiting from the self-attention mechanism, enable global modelling and have demonstrated impressive performance in surgical video analysis tasks such as surgical phase recognition [24, 25]. Nevertheless, the quadratic computational complexity of self-attention with respect to video sequence length significantly increases the computational burden, impeding its efficiency in handling long video scenarios.
@@
Figure 1: A surgical video from the SAR-RARP50 dataset, including errors of varying duration.
@@
Recently, state space models (SSMs) have proven highly effective in capturing long sequences in the natural language processing domain [26, 27]. Selective SSM [27], namely Mamba, distinguishes itself from other SSM-based methods by integrating a data-dependent SSM with a selection mechanism that employs parallel scanning, thereby achieving linear complexity while preserving global receptive fields. Inspired by the great success of Mamba, researchers have explored it in the computer vision domain, yielding promising performance. Zhu et al. [28] proposed a bidirectional Mamba (Vim) that scans images forward and backward to enhance spatial-aware understanding for image tasks, significantly improving computational efficiency while achieving higher performance compared to well-established Transformers. SegMamba [29] constructed a tri-oriented Mamba for 3D medical image segmentation. However, as depicted in Fig. 1, surgical videos contain errors with varied duration, necessitating the algorithm to consider temporal information at different granularities. Prior selective SSM-based works rarely consider modelling multi-scale temporal information, we thus showcase the effectiveness of a novel selective SSM-based method named SEDMamba, which captures temporal information across various scales to detect surgical errors in suturing from robotic-assisted radical prostatectomy (RARP) cases.
@@
    We propose a novel hierarchical model, SEDMamba, constructed with bottleneck multi-scale state space (BMSS) blocks to detect surgical errors in long videos from real RAS cases. To handle errors with variable duration, we design a fine-to-coarse temporal fusion (FCTF) using different dilated 1D convolutional layers to merge temporal information across different scale ranges. Besides, to further reduce computational complexity, we incorporate a bottleneck mechanism that compresses the spatial dimension of features, enabling selective SSM to concentrate on modelling the temporal dimension. To the best of our knowledge, this is the first attempt to exploit selective SSM in SED.
@@
    We present the first-of-its-kind, frame-level surgical error annotations of an in-vivo RAS dataset and apply them to train and validate the proposed SEDMamba model. Our work establishes a roadmap of how to utilize this new dataset for developing SED benchmarks.
@@
    The proposed SEDMamba outperforms state-of-the-art methods by 1.82-13.06% AUC and 3.8-18.86% AP while requiring 11.71-91.67% fewer model parameters and 40.71-92.32% fewer FLOPs.
@@ II-A Dataset Description and Annotation
A total of 48 videos, sampled at 60 Hz, from the open-source, SAR-RARP50 [30, 31], dataset of RARP cases were independently and blindly reviewed by two assessors. SAR-RARP50 focuses on the suturing of the dorsal venous complex (DVC) task and contains open-source annotations of eight gestures used as a task analysis sheet [30]. We developed a customized tool for the annotation of both procedural and executional errors using a previously validated suturing checklist and the OCHRA methodology [32]. This was summarized as a list of possible procedural and executional errors within any given suturing task, as seen in Table I. The duration of these videos varies from one minute to over eleven minutes, including the complete suturing process. Two assessors (one senior and one junior RAS surgeon with expertise in RAS education) analyzed the untrimmed videos sequentially frame-by-frame and annotated the start and end frame indexes of all errors. Fig. 2 illustrates three types of errors from our annotations. Inter-rater reliability was verified with Pearson’s correlation showing very strong agreement ($r$=0.8726, $p<$0.001) between the two assessors. Bland-Altman analysis shows average of 46.43 errors/case, and average difference of 5.87 errors between raters. Consensus was established for all disagreements via joint review.
Overall, 2507 error events were annotated, with E24-“Incorrect/poor instrument control” occurring the most frequently (686), followed by E1-“Multiple attempts” (510). The frequency of error types had statistically significant difference (Chi-squared test: $\chi 2$ = 1945.15, $p<$ 2.2e-16). The duration of errors varies, with the longest average duration being E19-“Suture entanglement”, averaging around 599 frames per case, while the shortest is E11-“Grasped at needle tip”, averaging 8 frames per case. The 48 videos comprise a total of 976037 frames. In the SED task, the frame is categorized as ‘normal’ if no error exists or ‘error’ if any type of error exists, resulting in 320355 error frames. To detect errors of varying duration, irrespective of their type, it is crucial to model temporal information across multiple scales.
@@
Multiple attempts (end of first attempt
@@
E7 Incorrect position along needle (normally second/third) 89 1737
@@ II-B Surgical Error Detection Model (SEDMamba)
This section introduces the preliminary concepts related to SEDMamba, including state space modelling and discretization. We then formulate the SEDMamba architecture and its core bottleneck multi-scale state space (BMSS) block.
@@ II-B2 Overview of SEDMamba
SEDMamba has a hierarchical architecture, shown in Fig. 3, comprising three BMSS blocks. Taking a video $V\in\mathbb{R}^{L\times C\times H\times W}$ with $L$ length, we first extract the spatial embedding sequence $\mathcal{E}\in\mathbb{R}^{L\times D}$ from a fixed DINOv2 [33], where $D$ is the spatial dimension. DINOv2 is a strong vision foundation model proven to extract task-agnostic visual spatial embedding and excel in video-based tasks [33]. Then, SEDMamba compresses $\mathcal{E}$ in the spatial dimension and captures fine-to-coarse temporal feature through each BMSS block, resulting in $i$-th hierarchical spatial-temporal representations with size $L\times\frac{D}{2^{i}}$. The 1D convolutional layer with 1 kernel size and 1 channel is then used for binary classification and outputs the final error probability prediction $P\in\mathbb{R}^{L\times 1}$. Note that each value in $P$ indicates the likelihood that the corresponding frame is erroneous. We use binary cross-entropy loss $\mathcal{L}_{bce}$ for supervised training.
@@ II-B3 Bottleneck Multi-scale State Space (BMSS) Block
II-B4 Fine-to-Coarse Temporal Fusion (FCTF)
To tackle varying duration of error and restricted receptive fields of convolutional layers, we propose the FCTF to capture and merge multi-scale temporal information. FCTF stacks three dilated 1D convolution layers with dilation rates of 2, 4, and 8 from top to bottom, enabling the fine-to-coarse temporal feature extraction scale by scale. Besides, stacking dilated convolution increases the receptive field without the need to increase the parameter number by increasing the kernel size. The receptive field expands exponentially as the number of stacked layers increases. Therefore, with a few parameters, we achieve a significantly large receptive field in the temporal dimension, which mitigates model overfitting and facilitates efficient error detection. The receptive field at each layer, with a kernel size of 3, can be calculated with the following formula:
$\text{ReceptiveField}(l)=2^{l+2}-1$ (5)
where $l\in[1,3]$ represents the layer number.
@@ II-B4 Fine-to-Coarse Temporal Fusion (FCTF)
20:                  $\left.\left.\text{Broadcast}^{G}(\text{Linear}_{\Delta}(x))\right)\right)$
@@ III-A Experimental Settings
All experiments are implemented in PyTorch on a RTX 4090 GPU. We use AdamW optimizer with initial learning rate of 1e-4 and maximum training epoch set at 200. We sampled SAR-RARP videos at 5Hz to get 40 videos with 46,586 normal frames and 18,540 error frames as the training set, and 8 videos with 7549 normal frames and 2096 error frames as the testing set. We use DINOv2 with fixed weights pretrained on ImageNet [33] as the visual extractor to generate spatial embedding sequences $\mathcal{E}$ from the video with $L$ frames as the model input. The compression factor $G$ in the BMSS block is set as 64. The channel $E$ of dilated 1D convolutional layers in FCTF is fixed as $\frac{D}{8}$. Following  [27], the state dimension $N$ of selective SSM is set to $16$.
III-B Evaluation Metrics
This work focuses on efficiently detecting errors in long surgical videos, formulating SED as a frame-level binary classification task. We therefore adopt the standard and widely used evaluation metrics–area under the curve (AUC) and average precision (AP)–at the frame level to assess performance. AUC and AP are both comprehensive metrics to measure the ability of a classifier to distinguish between positive (‘error’) and negative (‘normal’) classes. AUC refers to the area under the receiver operating characteristic (ROC) curve, which plots the true positive rate (TPR) against the false positive rate (FPR) across various threshold settings. Given a set of points on the ROC curve, where the points are ordered by increasing FPR, the AUC can be calculated using the trapezoidal rule:
$\mathrm{AUC}=\sum_{i=1}^{n-1}\left(\frac{TPR_{i+1}+TPR_{i}}{2}\right)\cdot\left(FPR_{i+1}-FPR_{i}\right)$ (6)
where $TPR_{i}=\frac{TP_{i}}{TP_{i}+FN_{i}}$ and $FPR_{i}=\frac{FP_{i}}{FP_{i}+TN_{i}}$ are the TPR and FPR at point $i$, respectively. TP, TN, FP, and FN represent the number of true positives, true negatives, false positives, and false negatives, respectively. $n$ is the number of thresholds.
AP is calculated as the sum of the precision values weighted by the change in recall:
$\mathrm{AP}=\sum_{i=1}^{n}\left(R_{i}-R_{i-1}\right)\cdot P_{i}$ (7)
where $R_{i}=\frac{TP_{i}}{TP_{i}+FN_{i}}$ and $P_{i}=\frac{TP_{i}}{TP_{i}+FP_{i}}$ are the recall and precision at the $i$-th threshold. Larger AUC and AP values indicate better performance. We report the mean and standard deviation from three runs with different random seeds.